OpenStudy (anonymous):

Find the area of the region enclosed by the curves and the y axis y=e^x+5 y=9

OpenStudy (anonymous):

ooh and the \(y\) axis!!!

OpenStudy (anonymous):

first need limits of integration \(y\) axis is \(x=0\) then solve \[e^x+5=9\] \[e^x=4\] \[x=\ln(4)\]

OpenStudy (anonymous):

yeah wich gave me 0 and 1.2 but i still couldnt find the right answer ;/

OpenStudy (anonymous):

or i was guessing that the limits were 0 and 1.2.. i am probably not right though

OpenStudy (anonymous):

\[\int _0^{\ln(4)}(9-e^x+5)dx\] \[\int_0^{\ln(4)}(4-e^x)dx\] etc

OpenStudy (anonymous):

gotta run. twilight zone you can do this integral mimi will help

OpenStudy (anonymous):

yaaayyyy that was right thank you so much!